# Learning to be Indifferent in Complex Decisions:

### v1

@@ 1 Introduction
We consider a standard decision problem under uncertainty where different alternatives may be available depending on the state of the world. However, we depart from the Bayesian paradigm by assuming that our decision-maker employs a simple learning heuristic to guide her strategy and its adjustments over time as she accumulates experience. Specifically, our decision-maker starts with initial assessments of the values of her alternatives, makes her choice by employing a mixed strategy that assigns higher probabilities to alternatives with higher assessments, and updates her assessments of the chosen alternatives in the direction of the observed payoff. Our learning model broadly belongs to the tradition of model-free reinforcement learning (RL), to the extent that the dynamics are driven solely by the observed payoffs. This particular genre of reinforcement learning with bandit feedback has been referred to as the Payoff-Assessment Learning model in the literature (Sarin and Vahid,, 1999), although we incorporate a noisy (smooth) version of it that uses a logit choice rule based on the assessments, similar to the one used in Cominetti et al., (2010). [^1] The logit (softmax) formulation has been viewed by some researchers (Rustichini et al.,, 2023) as providing a heuristic approach to model how a decision-maker could handle the exploitation/exploration trade-off, although the learning heuristic considered here is not forward-looking per se. In our study, we treat the parameter that governs the logit formulation as exogenous and focus on its limit, where the decision-maker almost surely picks the alternative(s) with the highest assessment(s).
@@
We believe that our modification of the usual payoff-assessment learning model is particularly relevant when the grand set of alternatives is too large to be considered extensively. In such cases, it seems highly plausible that the decision-maker would use coarse categories and implement the assessment device at the level of such categories rather than at the level of individual alternatives. This perspective in terms of categorical reasoning in the face of choice overload is well known in the psychological literature (Rosch and Lloyd,, 1978); however, to the best of our knowledge, it has not been explored in mathematical analyses of learning models. This is the main focus of our paper. It should be emphasized from the outset that, in our approach, the categories (referred to as similarity classes) are exogenously defined. One possible interpretation is that alternatives are characterized by vectors of attributes and the agent focuses primarily on those attributes that are salient while ignoring the others (Tversky,, 1972; Bordalo et al.,, 2012, 2013). From this perspective, a given similarity class corresponds to a specific profile of realizations of the salient attributes, potentially encompassing a large subset of alternatives. [^3] An alternative perspective is that a third party collects the data and organizes it into predefined categories. While this view aligns with the experimental framework in Jehiel and Singh, (2021), it is less relevant to our setting, which focuses on a single learning agent. Other potential mechanisms for aggregation could include limited memory or prevailing narratives.
@@
Our main contributions focus on analyzing the convergence properties of our learning model, particularly in the high-sensitivity limit. We start with the case of decision-makers equipped with at most two similarity classes where we provide a complete characterization - establishing convergence of the learning dynamics to the set of steady-states in the high-sensitivity limit. Specifically, we demonstrate that there always exists a steady-state that is locally asymptotically stable under CPAL dynamics, meaning that if initial valuations are close to those at the steady-state, the learning dynamics will converge to it. Furthermore, if the steady-state is unique, whether pure or mixed, it is globally asymptotically stable, meaning the learning dynamics will converge to it regardless of the initial conditions.
We then consider the general case of decision-makers with an arbitrary number of similarity classes and prove the following results. We verify that if a strict Valuation Equilibrium (VE) employing pure strategies exists, the steady-state that arises in its vicinity for a sufficiently large sensitivity parameter is locally asymptotically stable in the CPAL dynamics.
Our leading convergence results are characterized by varying the payoffs associated with trivial decision problems where all available alternatives belong to the same similarity class. When such payoffs are sufficiently high, [^4] Considering the additional cognitive cost associated with non-trivial choices (choice overload), this assumption is especially germane when differences in material payoffs are small, with the cognitive cost outweighing these differences. We believe this in particular applies to most experiments on complex decision-making. we demonstrate the emergence of a unique mixed steady-state, whose limit as the sensitivity parameter approaches infinity involves an equalization of assessments across at least two similarity classes. We establish that this unique mixed steady-state, whose limit features indifference(s) and a refinement of mixed VE, is globally asymptotically stable in the CPAL dynamics. By contrast, when trivial decision problems yield sufficiently low payoffs [^5] Our results do not address the cases where the payoffs attached to the trivial decision problems are at intermediate levels. We leave the exploration of these scenarios for future research., we show the existence of multiple steady-states, with at least one corresponding to a strict pure valuation equilibrium in the high-sensitivity limit. This steady-state is locally asymptotically stable for a sufficiently large sensitivity parameter, as established by our general local stability result for strict pure steady-states.
@@ 1.1 Literature Review
Our paper relates to various branches of literature. In the literature on learning, there is a long-standing practice of studying the stability properties of rest-points, beginning with the literature on fictitious play (Brown,, 1951), where Shapley, (1964) provided an early example of non-convergence. Convergence results in this domain were obtained for 2$\times$2 games (Brown,, 1951), two-player zero-sum games (Robinson,, 1951), dominance solvable games (Nachbar,, 1990) and potential games (Monderer and Shapley, 1996b, ; Monderer and Shapley, 1996a, ). The classical fictitious play model was extended to allow for stochastic (smooth) best-responses using the same logit formulation that we use (Fudenberg and Kreps,, 1993; Fudenberg and Levine,, 1998; Hofbauer and Sandholm,, 2002). The continuous-time long-run approximation of such models, developed by Benaïm, (1999), is adapted to and utilized in our learning model. Local stability results are derived by linearizing the obtained differential equations around the rest-points and verifying whether the real parts of the eigenvalues of the corresponding Jacobian matrix are all negative. Global stability results are established by either constructing a strict Lyapunov function that decays along all non-constant trajectories of the learning dynamics or leveraging the properties of cooperative dynamical systems (Smith,, 1995).
@@
Regarding categorization and its equilibrium consequences, beyond the Valuation Equilibrium (Jehiel and Samet,, 2007), the Analogy-based Expectation Equilibrium (Jehiel,, 2005) is worth mentioning. In this game-theoretic setting, players lump several states together into coarse categories to form aggregated predictions about opponents’ behavior. [^7] It can be viewed as the fictitious play counterpart of the approach developed in Jehiel and Samet, (2007).
In Section 2, we present the setup and the learning model, providing a continuous-time approximation as well as demonstrating the formal link between steady-states of the CPAL model and Valuation Equilibrium. Section 3 provides an example with multiple steady-states and another with a unique steady-state whose limit as the decision-maker almost surely selects the similarity class(es) with the highest assessment(s) involves mixing amid indifference. In Sections 4 and 5, we present our main analytical results for scenarios where the payoffs associated with trivial decision problems in which all alternatives belong to the same similarity class, are either sufficiently high or sufficiently low. Section 6 concludes the paper, highlighting open questions and avenues for future research.
@@ 2.2 Setup
$\mathbf{p}(\omega)=\sum_{\mathbf{\psi\in\Psi(\omega)}}f(\mathbf{\psi)}.$ (1)
@@
$r$$\psi_{1}$$(3)$$apples$$(2)$$limes$$\frac{1}{3}$$\psi_{2}$$(3)$$oranges$$(1)$$limes$$\frac{1}{3}$$\psi_{3}$$(4)$$oranges$$(1)$$apples$$\frac{1}{3}$
@@ 2.3 Learning Dynamics
More precisely, Alice encounters a stage decision tree $\mathcal{T^{\prime}}$ repeated infinitely. At each stage $k\in\mathbb{N}\cup\{0\}$, nature presents Alice with a choice problem $\omega\in\Omega$ with probability $\mathbf{p}(\omega)$. The set of similarity classes available to Alice at node $\omega$ is denoted by $\mathcal{S}_{\omega}\subseteq\mathcal{S}$. We refer to this set as the set of her pure strategies at node $\omega$. $\pi_{\omega}:\mathcal{S}_{\omega}\rightarrow\mathbb{R}$ is a bounded, generic payoff function at node $\omega$. We maintain the same notation to refer to the multi-linear extension of payoffs to the set of her mixed strategies. $\mathbf{\Delta_{\omega}}$ denotes the set of mixed strategies (probability vectors) over $\mathcal{S_{\omega}}$. Alice makes her choice at node $\omega$ in period $k$ employing a mixed strategy $\delta_{\omega,k}=\sigma_{\omega,k}(\mathbf{v}_{k})\in\mathbf{\Delta}_{\omega}$. Here, $\mathbf{v}_{k}=(v_{k}^{s})^{s\in\mathcal{S}}\in\mathbb{R}^{\mathcal{S}}$ is a vector of valuations that reflects her assessment of the payoff potential associated with each similarity class at stage $k$. We use the logit random choice model to characterize the mapping from the space of valuations to the space of mixed strategies. [^11] The logit random choice model has been widely used in stochastic fictitious play models of learning (Fudenberg and Levine,, 1998; Fudenberg and Kreps,, 1993; Hofbauer and Sandholm,, 2002), QRE models in game theory (McKelvey and Palfrey,, 1995; Goeree et al.,, 2016), and in discrete choice models in the empirical literature (Hausman and McFadden,, 1984; Anderson et al.,, 1992). It’s also commonly seen in the literature on reinforcement learning and deep learning, where it’s typically referred to as the softmax function. Thus, given a valuation vector $\mathbf{v}_{k}$, the probability that Alice chooses an alternative in similarity class $s\in\mathcal{S}_{\omega}$ at node $\omega$ in stage $k$ is given by the real-analytic ($C^{\omega}$) function:
$\sigma^{s}_{\omega,k}(\mathbf{v}_{k})=\frac{\exp(\beta v^{s}_{k})}{\sum_{j\in\mathcal{S}_{\omega}}\exp(\beta v^{j}_{k})}$ (3)
@@
Here, $\alpha_{k}\in(0,1)$ is a sequence of averaging factors such that $\sum_{k}\alpha_{k}=\infty$ and $\sum_{k}\alpha_{k}^{2}<\infty$ (Kushner and Yin,, 2003). These conditions are satisfied by setting $\alpha_{k}=(k+1)^{-1}$, which ensures that the valuation $v_{k}^{s}$ becomes the simple average of all observed payoffs for similarity class $s$ up to time $k$. This approach effectively gives equal weight to each observed payoff in the valuation, causing Alice’s sensitivity to new observations to diminish over time, while ensuring that future observations still exert a non-negligible impact. While Alice chooses her strategy based on her current valuation $\mathbf{v}_{k}$, the resulting payoffs, and thus the random vector $\tilde{\mathbf{v}}_{k}$, are influenced by these valuations. Equations (3) and (4) together describe a non-homogeneous Markov process that captures how her valuations evolve over time. This process can be interpreted as Alice exploring various similarity classes to understand their potential rewards and adjusting her future choices exploiting what she’s learned. The transition from one valuation $\mathbf{v}_{k}$ to the next $\mathbf{v}_{k+1}$ involves several steps: starting with her initial valuation, Alice employs a mixed strategy that assigns higher probabilities to alternatives in similarity classes with higher valuations, makes her choice, observes her realized payoff, and then updates her valuation for the relevant similarity class in the direction of the observed payoff. This exercise is iterated indefinitely generating a discrete-time stochastic process.
@@
First, we observe that in light of the continuous-time mean dynamics derived above, Alice’s transformation of the decision tree from $\mathcal{T}$ to $\mathcal{T^{\prime}}$ by collapsing multiple nodes into a single node using her similarity partition is indeed without loss of generality. Second, we note that our choice of the logit rule to model the mixed strategy map leads to the expected payoff terms being non-polynomial and rules out any chance of spelling out explicit, algebraic solutions to the system of ODEs for $\beta>0$. Of course, the ODE system is explicitly solvable for the trivial case of $\beta=0$ where Alice uniformly randomizes among all available similarity classes at any node $\omega$ [^15] A trivial case where the ODE system is explicitly solvable, even for large $\beta$, occurs when Alice possesses the finest similarity partition, $\mathcal{S}=\mathcal{C}$. That is, she does not cluster her alternatives and instead forms a distinct valuation for each alternative. Consequently, the system of ODEs decouples and simplifies to a linear form with constant $g_{s}=\mu_{\psi}(s)$ terms. The solutions exhibit exponential decay in time and asymptotically approach the actual payoffs leading to nearly optimal choices in the long-run. This exercise illustrates that clustering alternatives into coarse similarity classes significantly impacts the learning dynamics. We present a general result on the existence of steady-state solutions of the ODE system in Eq. (6). A steady-state solution is defined as a stationary system of valuations $\mathbf{v}^{*}\in\mathbb{R}^{\mathcal{S}}$ such that $\dot{\mathbf{v}}=0$ when evaluated at $\mathbf{v}^{*}$. We denote the set of steady-state solutions of the system of ODEs in Eq. (6) by $\mathcal{V}$. Applying Brouwer’s fixed point theorem, we prove the following existence result in the Appendix.
@@ 2.4 Connections with Valuation Equilibrium
The steady-states of the CPAL model whose existence is guaranteed in Theorem 1 can be interpreted as smooth variants of Valuation Equilibria, first introduced in the context of multi-player extensive form games by (Jehiel and Samet,, 2007).
@@
We readily verify that any steady-state of the CPAL model is a Smooth Valuation Equilibrium (SVE), and vice versa. For a finite $\beta$, the SVE is fully-mixed, meaning the corresponding vector of valuations lies in the interior of the convex hull of the payoffs. That is, at any node, Alice selects each available similarity class with strictly positive probability, according to the logit choice rule. However, as the logit parameter grows without bound ($\beta\uparrow\infty$), Alice becomes highly sensitive to differences in valuations. In this limit, the logit choice rule almost surely selects alternative(s) in the similarity class(es) with the highest valuation(s), $s\in\arg\max_{s\in\mathcal{S}_{\omega}}v_{s}$, at any node $\omega\in\Omega$ [^16] An alternative interpretation of the logit heuristic is that it approximates a noisy (stochastic) choice strategy that Alice employs based on her perception of the payoff performance of each similarity class. Essentially, this approach allows Alice to make errors in choosing her optimal similarity class, but it penalizes these errors in proportion to their severity. Specifically, the penalty incurred for making a sub-optimal choice is exponentially proportional to the payoff loss associated with that choice. As $\beta$ increases indefinitely, the cost of these mistakes becomes prohibitively high, effectively driving Alice’s behavior toward the optimal strategy. This property is used to show that as $\beta\uparrow\infty$, the corresponding smooth valuation equilibria lie in an arbitrarily small neighborhood of some valuation equilibrium. The latter requires that in each state $\omega$, Alice chooses similarity classes $s\in\arg\max_{s\in\mathcal{S}_{\omega}}v_{s}$ and valuations are consistent in the sense that $v_{s}=\frac{\sum_{\omega\in\Omega:s\in\mathcal{S}_{\omega}}p(\omega)\sigma_{\omega}^{s}(v)\pi_{\omega}(s)}{\sum_{\omega\in\Omega:s\in\mathcal{S}_{\omega}}p(\omega)\sigma_{\omega}^{s}(v)}$.
@@
It’s worth mentioning that while each fixed point of the CPAL dynamics in the high-sensitivity limit is a valuation equilibrium, there exist valuation equilibria in certain decision trees that cannot be characterized as the limiting fixed points of the continuous-time CPAL dynamics as $\beta\uparrow\infty$. An example illustrating this point is provided in the Online Appendix. Consequently, the set of smooth valuation equilibria in the high-sensitivity limit ($\beta\uparrow\infty$) of the CPAL model offers a refinement of the set of VE in a decision tree.
@@
Our primary focus in this paper is on the asymptotic convergence properties of the CPAL dynamics. Eq. (6) defines a real, autonomous, smooth dynamical system on the space of valuations. At any time, $t\in\mathbb{R}$, the state of the dynamical system is given by a vector of valuations, $\mathbf{v}(t)\in\mathbb{R}^{\mathcal{S}}$. The rest-points of the dynamical system are elements $\mathbf{v}^{*}$ of the set $\mathcal{V}$ such that the time-derivative equals zero at these points. In the following sections, we investigate the asymptotic stability [^17] An equilibrium of a dynamical system is locally asymptotically stable if, for any initial condition sufficiently close to the equilibrium, the solution trajectory remains close (Lyapunov stability) and asymptotically converges to the equilibrium (attractivity). Global asymptotic stability refers to the property of an equilibrium where all solutions of the dynamical system, regardless of the initial conditions, asymptotically converge to the equilibrium. Refer to the Online Appendix for a formal treatment of asymptotic stability and the Linearization Theorem. of the rest-points. For local stability, we examine the effects of small perturbations on the long-run behavior of the CPAL model in neighborhoods of its rest-points. In particular, we linearize Eq. (6) around the rest-points and analyze the sign of the real parts of the eigenvalues of the corresponding Jacobian matrices to determine the local asymptotic stability of the CPAL model at its rest-points. For global stability results of the continuous-time process, we either construct a strict Lyapunov function that decays along all non-constant trajectories of the CPAL dynamics or leverage the convergence properties of monotone cooperative dynamical systems.
@@
Pemantle, (1990) shows that the discrete-time system in Eq.(4) has a probability zero of converging to a linearly unstable steady-state of the continuous-time process in Eq.(6), provided that there is a non-negligible amount of noise in the evolution of every component of the system. Benaïm, (1999) shows that every locally asymptotically stable steady-state of the continuous-time process has strictly positive probability of being the long-run outcome of the discrete-time process, again provided that there is non-negligible noise in the system. Benaïm, (1999) also shows that if the continuous-time process has a unique steady-state that is a global attractor, then it is the unique element of the internally chain-transitive set and the discrete-time process converges to it almost surely.

### v2

@@ 1 Introduction
We consider a standard decision problem under uncertainty where different alternatives may be available depending on the state of the world. However, we depart from the Bayesian paradigm by assuming that our decision-maker employs a simple learning heuristic to guide her strategy and its adjustments over time as she accumulates experience. Specifically, our decision-maker starts with initial assessments of the values of her alternatives, makes her choice by employing a mixed strategy that assigns higher probabilities to alternatives with higher assessments, and updates her assessments of the chosen alternatives in the direction of the observed payoff. Our learning scheme broadly belongs to the tradition of model-free reinforcement learning (RL), to the extent that the dynamics are driven solely by the observed payoffs. This particular genre of reinforcement learning with bandit feedback has been referred to as the Payoff-Assessment Learning model in the literature (Sarin and Vahid,, 1999), although we incorporate a noisy (smooth) version of it that uses a logit choice rule based on the assessments, similar to the one used in Cominetti et al., (2010). [^1] The logit (softmax) formulation has been viewed by some researchers (Rustichini et al.,, 2023) as providing a heuristic approach to model how a decision-maker could handle the exploitation/exploration trade-off, although the learning heuristic considered here is not forward-looking per se. In our study, we treat the parameter that governs the logit formulation as exogenous and focus on its limit where the decision-maker almost surely picks the alternative(s) with the highest assessment(s).
@@
We believe that our modification of the usual payoff-assessment learning model is particularly relevant when the grand set of alternatives is too large to be considered extensively. In such cases, it seems highly plausible that the decision-maker would use coarse categories and implement the assessment device at the level of such categories rather than at the level of individual alternatives. This perspective in terms of categorical reasoning in the face of choice overload is well known in the psychological literature (Rosch and Lloyd,, 1978); however, to the best of our knowledge, it has not been explored in mathematical analyses of learning models. This is the main focus of our paper. It should be emphasized from the outset that, in our approach, the categories (referred to as similarity classes) are exogenously defined. One possible interpretation is that alternatives are characterized by vectors of attributes and the agent focuses primarily on those attributes that are salient while ignoring the others (Tversky,, 1972; Bordalo et al.,, 2012, 2013). From this perspective, a given similarity class corresponds to a specific profile of realizations of the salient attributes, potentially encompassing a large subset of alternatives. [^3] An alternative perspective is that a third party collects the data and organizes it into predefined categories. While this view aligns with the experimental framework in Jehiel and Singh, (2021), it is less relevant to our setting, which focuses on a single learning agent. Other potential mechanisms for aggregation include limited memory, unlabeled data, or prevailing narratives.
@@
Our main contributions focus on analyzing the convergence properties of our learning model, particularly in the high-sensitivity limit. We start with the case of decision-makers equipped with at most two similarity classes where we provide a complete characterization - establishing convergence of the learning dynamics to the set of steady-states in the high-sensitivity limit. Specifically, we demonstrate that there always exists a steady-state that is locally asymptotically stable under the CPAL dynamics, meaning that if the initial assessments are close to those at the steady-state, the learning dynamics will converge to it. Furthermore, if the steady-state is unique, whether pure or mixed, it is globally asymptotically stable, meaning the learning dynamics will converge to it regardless of the initial conditions.
We then consider the general case of decision-makers with an arbitrary number of similarity classes and prove the following results. We verify that if a strict Valuation Equilibrium (VE) employing pure strategies exists, the steady-state that arises in its vicinity for a sufficiently large sensitivity parameter is locally asymptotically stable under the CPAL dynamics.
Our leading convergence results are characterized by varying the payoffs associated with trivial decision problems where all available alternatives belong to the same similarity class. When such payoffs are sufficiently high, [^4] Considering the additional cognitive cost associated with non-trivial choices (choice overload), this assumption is especially germane when differences in material payoffs are small, with the cognitive cost outweighing these differences. We believe this in particular applies to most lab experiments on complex decision-making. we demonstrate the emergence of a unique mixed steady-state, whose limit as the sensitivity parameter approaches infinity involves an equalization of assessments across at least two similarity classes. We establish that this unique mixed steady-state, whose limit features indifference(s) and a refinement of mixed VE, is globally asymptotically stable in the CPAL dynamics. By contrast, when trivial decision problems yield sufficiently low payoffs [^5] Our results on the stability of mixed equilibria do not address the cases where the payoffs attached to trivial decision problems are at intermediate levels. We leave the exploration of these scenarios for future research., we show the existence of multiple steady-states, with at least one corresponding to a strict pure valuation equilibrium in the high-sensitivity limit. This steady-state is locally asymptotically stable for a sufficiently large sensitivity parameter, as established by our general local stability result for strict pure steady-states.
@@ 1.1 Literature Review
[…]
@@
Regarding categorization and its equilibrium consequences, beyond the Valuation Equilibrium (Jehiel and Samet,, 2007), the Analogy-based Expectation Equilibrium (Jehiel,, 2005) is also worth mentioning. In this game-theoretic setting, players lump several states together into coarse categories to form aggregated beliefs about opponents’ behavior. [^7] It can be viewed as the fictitious play counterpart of the approach developed in Jehiel and Samet, (2007).
In Section 2 of the paper, we present the setup and the learning model, providing a continuous-time approximation as well as demonstrating the formal link between steady-states of the CPAL model and Valuation Equilibrium. Section 3 provides an example with multiple steady-states and another with a unique steady-state whose limit as the decision-maker almost surely selects the similarity class(es) with the highest assessment(s) involves mixing amid indifference. In Sections 4 and 5, we present our main analytical results for scenarios where the payoffs associated with trivial decision problems in which all alternatives belong to the same similarity class, are either sufficiently high or sufficiently low. Section 6 concludes the paper, highlighting open questions and avenues for future research.
@@ 2.2 Setup
$p(\omega)=\sum_{\mathbf{\psi\in\Psi(\omega)}}f(\mathbf{\psi)}.$ (1)
@@
$r$$\psi_{1}$$(3)$$apples$$(2)$$limes$$\frac{1}{3}$$\psi_{2}$$(3)$$lemons$$(1)$$limes$$\frac{1}{3}$$\psi_{3}$$(4)$$lemons$$(1)$$apples$$\frac{1}{3}$
@@ 2.3 Learning Dynamics
More precisely, Alice encounters a stage decision tree $\mathcal{T^{\prime}}$ repeated infinitely. At each stage $k\in\mathbb{N}\cup\{0\}$, nature presents Alice with a choice problem $\omega\in\Omega$ with probability $p(\omega)$. The set of similarity classes available to Alice at node $\omega$ is denoted by $\mathcal{S}_{\omega}\subseteq\mathcal{S}$. We refer to this set as the set of her pure strategies at node $\omega$. $\pi_{\omega}:\mathcal{S}_{\omega}\rightarrow\mathbb{R}$ is a bounded, generic payoff function at node $\omega$. We maintain the same notation to refer to the multi-linear extension of payoffs to the set of her mixed strategies. $\mathbf{\Delta_{\omega}}$ denotes the set of mixed strategies (probability vectors) over $\mathcal{S_{\omega}}$. Alice makes her choice at node $\omega$ in period $k$ employing a mixed strategy $\delta_{\omega,k}=\sigma_{\omega,k}(\mathbf{v}_{k})\in\mathbf{\Delta}_{\omega}$. Here, $\mathbf{v}_{k}=(v_{k}^{s})^{s\in\mathcal{S}}\in\mathbb{R}^{\mathcal{S}}$ is a vector of valuations that reflects her assessment of the payoff potential associated with each similarity class at stage $k$. We use the logit stochastic choice model to characterize the mapping from the space of valuations to the space of mixed strategies. [^11] The logit choice model has been widely used in stochastic fictitious play models of learning (Fudenberg and Levine,, 1998; Fudenberg and Kreps,, 1993; Hofbauer and Sandholm,, 2002), QRE models in game theory (McKelvey and Palfrey,, 1995; Goeree et al.,, 2016), and in discrete choice models in the empirical literature (Hausman and McFadden,, 1984; Anderson et al.,, 1992). It’s also commonly seen in the literature on reinforcement learning and deep learning, where it’s typically referred to as the softmax function. Thus, given a valuation vector $\mathbf{v}_{k}$, the probability that Alice chooses an alternative in similarity class $s\in\mathcal{S}_{\omega}$ at node $\omega$ in stage $k$ is given by the real-analytic ($C^{\omega}$) function:
$\sigma^{s}_{\omega,k}(\mathbf{v}_{k})=\frac{\exp(\beta v^{s}_{k})}{\sum_{j\in\mathcal{S}_{\omega}}\exp(\beta v^{j}_{k})},$ (3)
@@
Here, $\alpha_{k}\in(0,1)$ is a sequence of averaging factors such that $\sum_{k}\alpha_{k}=\infty$ and $\sum_{k}\alpha_{k}^{2}<\infty$ (Kushner and Yin,, 2003). These conditions are satisfied, for e.g., by setting $\alpha_{k}=(k+1)^{-1}$, meaning that the valuation $v_{k}^{s}$ becomes the simple average of all observed payoffs for similarity class $s$ up to time $k$. This approach effectively gives equal weight to each observed payoff in the valuation, causing Alice’s sensitivity to new observations to diminish over time, while ensuring that future observations still exert a non-negligible impact. While Alice chooses her strategy based on her current valuation $\mathbf{v}_{k}$, the resulting payoffs, and thus the random vector $\tilde{\mathbf{v}}_{k}$, are influenced by these valuations. Equations (3) and (4) together describe a non-homogeneous Markov process that captures how her valuations evolve over time. This process can be interpreted as Alice exploring various similarity classes to understand their potential rewards and adjusting her future choices exploiting what she has learned. The transition from one valuation $\mathbf{v}_{k}$ to the next $\mathbf{v}_{k+1}$ involves several steps: starting with her initial valuation, Alice employs a mixed strategy that assigns higher probabilities to alternatives in similarity classes with higher valuations, makes her choice, observes her realized payoff, and then updates her valuation for the relevant similarity class in the direction of the observed payoff. This exercise is iterated indefinitely generating a discrete-time stochastic process.
@@
First, we observe that in light of the continuous-time mean dynamics derived above, our transformation of the decision tree from $\mathcal{T}$ to $\mathcal{T^{\prime}}$ by collapsing multiple analogous nodes into a single node using Alice’s similarity partition is indeed without loss of generality. Second, we note that our choice of the logit rule to model the mixed strategy map leads to the expected payoff terms being non-polynomial and rules out any chance of spelling out explicit, algebraic solutions to the system of ODEs for $\beta>0$. Of course, the ODE system is explicitly solvable for the trivial case of $\beta=0$ where Alice uniformly randomizes among all available similarity classes at any node $\omega$ [^15] A trivial case where the ODE system is explicitly solvable, even for large $\beta$, occurs when Alice possesses the finest similarity partition, $\mathcal{S}=\mathcal{C}$. That is, she does not cluster her alternatives and instead forms a distinct valuation for each alternative. Consequently, the system of ODEs decouples and simplifies to a linear form with constant $g_{s}=\mu_{\psi}(s)$ terms. The solutions exhibit exponential decay in time and asymptotically approach the actual payoffs leading to nearly optimal choices in the long-run. This exercise illustrates that clustering alternatives into coarse similarity classes significantly impacts the learning dynamics. We present a general result on the existence of steady-state solutions of the ODE system in Eq. (6). A steady-state solution is defined as a stationary system of valuations $\mathbf{v}^{*}\in\mathbb{R}^{\mathcal{S}}$ such that $\dot{\mathbf{v}}=0$ when evaluated at $\mathbf{v}^{*}$. We denote the set of steady-state solutions of the system of ODEs in Eq. (6) by $\mathcal{V}$. Applying Brouwer’s fixed point theorem, we prove the following existence result in the Appendix.
@@ 2.4 Connections with Valuation Equilibrium
The steady-states of the continuous-time CPAL model whose existence is guaranteed in Theorem 1 can be interpreted as smooth variants of Valuation Equilibria, first introduced in the context of multi-agent extensive form games by Jehiel and Samet, (2007).
@@
We readily verify that any steady-state of the CPAL model is a Smooth Valuation Equilibrium (SVE), and vice versa. For a finite $\beta$, any SVE is fully-mixed, meaning the corresponding vector of valuations lies in the interior of the convex hull of the payoffs. That is, at any node, Alice selects each available similarity class with strictly positive probability, according to the logit choice rule. However, as the logit parameter grows without bound ($\beta\uparrow\infty$), Alice becomes highly sensitive to differences in valuations. In this limit, the logit choice rule almost surely selects the similarity class(es) with the highest valuation(s), $s\in\arg\max_{s\in\mathcal{S}_{\omega}}v_{s}$, at any node $\omega\in\Omega$ [^16] An alternative interpretation of the logit heuristic is that it approximates a noisy (stochastic) choice strategy that Alice employs based on her perceived payoff performance of each similarity class. While this approach allows Alice to make errors in choosing her optimal similarity class, it penalizes these errors in proportion to their severity. Specifically, the penalty incurred for making a sub-optimal choice is exponentially proportional to the payoff loss associated with that choice. As $\beta$ increases indefinitely, the cost of these mistakes becomes prohibitively high, driving Alice’s behavior toward the optimal strategy. This property is used to show that as $\beta\uparrow\infty$, the corresponding smooth valuation equilibria lie in an arbitrarily small neighborhood of some valuation equilibrium. The latter requires that in each state $\omega$, Alice chooses similarity class(es) $s\in\arg\max_{s\in\mathcal{S}_{\omega}}v_{s}$ and her valuations are consistent in the sense that $v_{s}=\frac{\sum_{\omega\in\Omega:s\in\mathcal{S}_{\omega}}p(\omega)\sigma_{\omega}^{s}(v)\pi_{\omega}(s)}{\sum_{\omega\in\Omega:s\in\mathcal{S}_{\omega}}p(\omega)\sigma_{\omega}^{s}(v)}$.
@@
It’s worth mentioning that while each fixed point of the CPAL dynamics in the high-sensitivity limit is a valuation equilibrium, there exist valuation equilibria in certain decision trees that cannot be characterized as the limiting fixed points of the continuous-time CPAL dynamics as $\beta\uparrow\infty$. An example illustrating this point is provided in Sec. 3.1 of the Online Appendix. Consequently, the set of smooth valuation equilibria in the high-sensitivity limit ($\beta\uparrow\infty$) of the CPAL model offers a refinement of the set of VE in a decision tree.
@@
Our primary focus in this paper is on the asymptotic convergence properties of the CPAL dynamics. Eq. (6) defines a real, autonomous, smooth dynamical system on the space of valuations. At any time, $t\in\mathbb{R}$, the state of the dynamical system is given by a vector of valuations, $\mathbf{v}(t)\in\mathbb{R}^{\mathcal{S}}$. The rest-points of the dynamical system are elements $\mathbf{v}^{*}$ of the set $\mathcal{V}$ such that the time-derivative equals zero at these points. In the following sections, we investigate the asymptotic stability [^17] An equilibrium of a dynamical system is locally asymptotically stable if, for any initial condition sufficiently close to the equilibrium, the solution trajectory remains close (Lyapunov stability) and asymptotically converges to the equilibrium (attractivity). Global asymptotic stability refers to the property of an equilibrium where all solutions of the dynamical system, regardless of the initial conditions, asymptotically converge to the equilibrium. Refer to Sec. 4.1. of the Online Appendix for a formal treatment of asymptotic stability and the Hartman-Grobman (Linearization) Theorem. of the rest-points. For local stability, we examine the effects of small perturbations on the long-run behavior of the CPAL model in neighborhoods of its rest-points. In particular, we linearize Eq. (6) around the rest-points and analyze the sign of the real parts of the eigenvalues of the corresponding Jacobian matrices to determine the local asymptotic stability of the CPAL model at its rest-points. For global stability results of the continuous-time process, we either construct a strict Lyapunov function that decays along all non-constant trajectories of the CPAL dynamics or leverage the convergence properties of monotone cooperative dynamical systems (Smith,, 1995).
@@
Pemantle, (1990) shows that a discrete-time stochastic system as in Eq. (4) has a probability zero of converging to a linearly unstable steady-state of the continuous-time process in Eq. (6), provided that there is a non-negligible amount of noise in the evolution of every component of the system. Benaïm, (1999) shows that every locally asymptotically stable steady-state of the continuous-time process has strictly positive probability of being the long-run outcome of the discrete-time process, again provided that there is non-negligible noise in the system. Benaïm, (1999) also shows that if the continuous-time process has a unique steady-state that is a global attractor, then it is the unique element of the internally chain-transitive set and the discrete-time process converges to it almost surely. In light of these results, we focus the remainder of the paper on analyzing the convergence properties of the continuous-time CPAL dynamics in Eq. (6).